OpenStudy (anonymous):

x-(average of 6 numbers) = 6,-4,6,2,2,0 there are 6 x's, but i need to find the average of the 6 numbers given the equation above. How do i do that?

OpenStudy (anonymous):

so is it : x - (average of 6,-4,6,2,2,0) ? i don't get the equation...

OpenStudy (anonymous):

no, x- (average) = 6 y - (average) = -4 ect.... there are 6 variables who's average are what im looking for so basically, (variable)-(average of variables)=6,-4,6,2,2,0 where there are 6 variable for 6 different answers

OpenStudy (anonymous):

ok sorry... i m as clueless bout this stuff... haven't learnt it yet... trying asking someone chatting

OpenStudy (anonymous):

how should i word this question to help not confuse people?

OpenStudy (anonymous):

ok i will be here for a few sec so gotta be quick... r u asking for average of 6? (eg)

OpenStudy (anonymous):

i don't get the 6y part

OpenStudy (anonymous):

see if you can follow, like, there are 6 different variable, when added up, and then when each individual variable is subtracted from the average of all of them, the answers are given, 6 different answers, some variable may be used 2 times, but in the end, there are 6 numbers averaged and subtracted from each of the individual numbers to get the end results

OpenStudy (anonymous):

idk if i just helped you or got you more confused, but thats the best i can explain it

OpenStudy (anonymous):

ok... i m abit clear but still a bit confused.. i will ask in the chat box... sorry:( couldn't help

OpenStudy (anonymous):

ok just ask ur teacher, she/he will explain it clearer:)

OpenStudy (anonymous):

ok, thanks anyway though!

OpenStudy (anonymous):

thanks for the medal!:D sorry for not being able to ans though... hope u will get the answer soon!

OpenStudy (anonymous):

thank you! (:

OpenStudy (anonymous):

i found out and feel rly stupid now

OpenStudy (anonymous):

i just found out and feel rly stupid

OpenStudy (anonymous):

u found out? yay!:P don't feel stupid... as long as u found it out then good!

OpenStudy (anonymous):

you can pick any number to be the average and just follow the equation to make it fit, pick any number as the average, add it to the numbers and in the end the equation workd

OpenStudy (anonymous):

haha... i m still clueless... but u got it! that's wat matters rite?:P

OpenStudy (anonymous):

g2g, bye

OpenStudy (anonymous):

bye! have a great day ahead!:D